# CLIPLoss and Norm-Based Data Selection Methods for Multimodal Contrastive Learning

### v1

@@ 4 Experimental Results
In this section, we evaluate the performance of negCLIPLoss and NormSim, aiming to address the following questions: Q1: Given a fixed CLIP teacher model, can our methods more effectively utilize CLIP embeddings for data filtering? Q2: Are our methods applicable to diverse CLIP teacher models with varying architectures or different pretrained datasets? Q3: How does our method compare to other leading approaches that utilize external models or multimodal datasets? Additionally, could our method be compatible with these methods and enhance their effectiveness?
@@ 4.1 Setup
We adhere to the standardized training and evaluation protocols of the DataComp benchmark [1]. Training configuration. We employ the medium-scale training configuration of DataComp (DataComp-medium). It provides a substantial dataset comprising 128 million low-quality, web-curated image-text pairs to be filtered. Once the data subset is obtained by some data filtering strategy, it will be used to train a fixed CLIP-B/32 model in a fixed training budget that allows the model to pass 128 million data points an epoch. Therefore, smaller subsets will be repeated more frequently, ensuring a fair comparison. We note that the size of the DataComp dataset becomes smaller over time since some URLs of images become invalid [^3] See https://github.com/mlfoundations/datacomp/issues/3. Similar issues are proposed by $\mathbb{D}^{2}$ pruning [14]., and we only successfully downloaded about 110M data. Therefore, the results of baselines on the leaderboard do not apply to our datasets, and we reproduce all the top baselines on the leaderboard with their public UIDs of the selected data.
Evaluation. We measured the model performance on 38 downstream datasets including image classification and retrieval tasks followed by DataComp. The image classification tasks contain ImageNet-1k [23], ImageNet distribution shifts [24, 25, 26, 27], 11 datasets from the Visual Task Adaptation Benchmark (VTAB) [28] and 3 datasets from WILDS [29, 30]. Retrieval datasets contain Flickr30k [31], MSCOCO [32] and WinoGAViL [33].
Teacher model architecture. Our experiments utilize two architectures for OpenAI’s CLIP teacher models: ViT-L/14 and ViT-B/32. Additionally, we use the public version of DFN (DFN-P) proposed by Fang et al. [2] as a teacher model, and its architecture is also ViT-B/32.
@@ 4.2 Baselines
We restate the three current research directions mentioned before based on how much external resources are employed: (D1) using OAI CLIP alone while optimizing embedding employment strategies, (D2) training and using a more advanced CLIP embedding model based on external data, and (D3) utilizing non-CLIP external models to aid data selection. It is important to note that D2 and D3 may also incorporate strategies from D1. For example, CLIPScore (D1) has been used in almost all the top methods. Therefore, we categorize baselines by the largest possible category they encompass. According to the above categorization, we summarize the baselines we used in our experiments as follows. Please refer to Fig. 4 and Appendix C.4 for more details.
D1: OAI CLIP embedding only. The learner can only access the pretraining dataset (like DataComp-medium), the original OAI CLIP teacher model that is used to extract embeddings, and some target data of the downstream tasks which is much smaller than the pretraining dataset (like ImageNet-1k). In this category, we don’t use any existing external non-CLIP models or any newly trained CLIP model based on external multi-modal dataset. In detail, This category includes (1) CLIPScore [34], which only uses CLIPScore for filtering as we mentioned before. (2) Image-based filtering [1], which uses ImageNet-1K training data as the downstream target data for data filtering. It applies k-means clustering to the image embeddings of training data and selects clusters closest to the ImageNet-1K embeddings. Gadre et al. [1] also try to combine image-based filtering and CLIPScore together. (3) $\mathbb{D}^{2}$ Pruning [14], which represents the dataset as an undirected graph and selects the data by combining difficulty and diversity. They use the CLIP score to initialize their graph.
D2, D3: Accessible external model and multi-modal data. All the current top baselines enable the learner to utilize external resources, either to train a better CLIP teacher model or to help filtering using existing models’ properies. In detail, (1) DFN [2] trains another CLIP data filtering network via external high-quality data. Their currently public model (DFN-P) is trained on CC12M [35] + CC3M [36] + SS15M [37], while the best DFN is trained on nonpublic HQITP-350M [2], which is even larger than DataComp-medium. (2) HYPE [3] leverages hyperbolic embeddings (different from CLIP embedding) and the concept of entailment cones to filter out samples with meaningless or underspecified semantics, enhancing the specificity of each sample. (3) HYPE $\cup$ DFN proposed by [3] samples subset separately for each method and then merge them. This is the state-of-the-art method on the DataComp benchmark for medium size. (4) Other methods including T-MARS [8], Devils [10], MLM [38], which leverage external models such as text detection model FAST [9], vision-language model (VLM) BLIP-2 [12] and VLM LLaVA-1.5 [39, 40] to heuristically select the more helpful data, like non-text-only pictures or images with MNIST-style digits. See details in Appendix C.4.
Cross-setting comparison. We make these separations for fair comparison. Intuitively, performance should be ranked as D2, D3 > D1. However, our results show that cross-setting comparisons are possible and our D1 methods can perform similar or even better than most of D3 methods.
4.3 Main Results and Discussions
Filtering Strategy Dataset IN-1k IN Dist. Shift VTAB Retrieval Avg.
@@ 4.3 Main Results and Discussions
No filtering [1] 110M 17.3 15.0 25.2 21.3 25.6
CLIPScore (20%) [34] 22M 25.4 22.7 31.8 22.0 31.0
CLIPScore (30%) [34] 33M 26.4 23.6 32.6 24.5 32.2
Image-based [1] 24M 25.5 21.9 30.4 24.6 29.9
CLIPScore (30%) $\cap$ Image-based  [1] 11M 27.4 23.9 31.9 21.4 30.8
$\mathbb{D}^{2}$ Pruning [14] 22M 23.2 20.4 31.4 18.7 29.5
@@
4.3.1 Comparision on D1 Category (Q1)
In Table 1, we compare the D1 methods where only the OAI CLIP model is allowed to be used.
Our Methods leverage OAI CLIP-L/14 better. First, negCLIPLoss outperforms CLIPScore on all metrics, regardless of whether it is used alone or combined with other methods. These results support our claim that negCLIPLoss can more accurately estimate the data quality.
Table 2: Results on DataComp-medium from the top methods that use only OpenAI’s CLIP-B/32 model or public version of DFN (DFN-P). “$\text{NormSim}_{\infty}^{\text{B/32}}$” represents using OAI CLIP-B/32 to calculate $\text{NormSim}_{\infty}$.
@@ 4.3.1 Comparision on D1 Category (Q1)
[…]
4.3.2 Try Other Teacher Models (Q2)
To evaluate whether our method applies to other CLIP teacher models, we replaced OAI CLIP-L/14 with OAI CLIP-B/32 and DFN-P as embedding models. We compare the best baseline “CLIPScore” with our “negCLIPLoss” and best strategy “negCLIPLoss $\cap$ $\text{NormSim}_{\infty}$(Target)” as shown in Table 2 and Appendix D.2. Note that the original DFN paper selects a subset comprising 19.2M data points, which accounts for approximately $17.5\%$ of our dataset and $15\%$ of their dataset, we incorporate these sampling ratios into our comparison.
negCLIPLoss can be applied to different CLIP embedding models. Our proposed negCLIPLoss, as a replacement of CLIPScore, not only leads to better performance compared to all the other baselines using OAI CLIP-L/14 as shown in Table 1, but also achieves universal improvement on the other two CLIP embedding models, OAI CLIP-B/32 and DFN-P as shown in Table 2. Our methods can consistently outperform all downstream tasks for different filtering ratios and models, like a 0.5%-5.4% increase on ImageNet-1k.
Embedding required by NormSim should have good downstream performance. When combining negCLIPLoss with $\text{NormSim}_{\infty}$, OAI CLIP-B/32 and DFN-P exhibit completely different behaviors. The former obtains results even better than those in Table 1, which uses OAI CLIP-L/14 as the teacher model, while DFN-P achieves results even worse than using negCLIPLoss alone (see "negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)" and "negCLIPLoss (20%)/(30%)"). The reason is that, unlike OAI CLIP-B/32, DFN-P is specially designed for data filtering at the expense of downstream task performance, as claimed by its authors. For example, the ImageNet-1k accuracy for DFN-P, OAI CLIP-B/32, and OAI CLIP-L/14 are 45%, 63%, and 75%, respectively. This indicates that the embeddings obtained from DFN on target data might be highly unreliable, leading to inaccurate similarity calculations between training and target data. To support this, if we use DFN-P to evaluate negCLIPLoss but utilize OAI CLIP-B/32 for calculating NormSim, as shown in "negCLIPLoss (17.5%) $\cap$ $\text{NormSim}_{\infty}^{\text{B/32}}$(Target)", we can further improve the results compared to using negCLIPLoss alone. Its average performance on 38 tasks is even higher than utilizing the best DFN (trained on HQITP-350M) with CLIPScore, as shown in Table 3.
@@ 4.3.3 Comparison with D2 & D3 Categories (Q3)
[…]
Table 3: Results of all D1&D2&D3 top methods on DataComp-medium. The results of MLM [38] are from their paper, while all other baselines are reproduced on our downloaded dataset using their official UIDs. “Ours (20%)” refers to use “negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)” to get 20% of original data, while “Ours (10%)” denotes applying “negCLIPLoss (20%) $\cap$ $\text{NormSim}_{\infty}$(Target)” to get 10%. And we use “*” to indicate the case where we choose the intersection of the data selected by using OAI CLIP-B/32 and OAI CLIP-L/14 separately, which results in about 15M data for “Ours (20%)*” and 7.4M data for “Ours (10%)*”.
Type Filtering Strategy Dataset IN-1k IN Dist. Shift VTAB Retrieval Avg.
Size (1) (5) (11) (3) (38)
D2 DFN [2] 16M 36.0 30.1 36.2 27.0 35.4
D3 T-MARS [8] 22M 30.8 26.3 34.8 25.4 34.1
D3 Devil [10] 20M 31.0 26.7 35.9 24.7 34.5
D3 MLM [38] 38M 30.3 25.6 36.0 29.0 34.5
D3 HYPE [3] 10M 30.3 25.8 34.3 22.2 31.9
D3 DFN $\cup$ HYPE [3] 20M 36.4 30.8 38.5 28.0 36.8
@@
Our methods can outperform most of the D3 methods. In Table 3, we show that without using any external models or data, our best combination, i.e., using OAI CLIP-B/32 for “negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)” (Ours (20%)), still outperforms all methods except DFN and “DFN $\cup$ HYPE”. This answers the first part of Q3 and further indicates that some external models may be redundant since CLIP embeddings already contain necessary information.
We can further improve the SOTA method. We show that our model can further boost the performance of the current SOTA method “HYPE $\cup$ DFN” by 0.9% on both ImageNet-1k and on average 38 downstream tasks, and close results can be achieved even without combining HYPE which utilizes the external embedding model MERU [41]. Here we use the data selected by both OAI CLIP-B/32 and L/14, which we found is more robust than using one of them alone. Our better results answer the second part of Q3, that is, our methods can be compatible with other D2&D3 methods.
5 Conclusion and Limitation
In this paper, we introduce two metrics, negCLIPLoss and NormSim, to enhance data selection in multimodal contrastive learning without relying on external resources. negCLIPLoss provides a more accurate quality metric compared to the commonly used CLIPScore, while NormSim measures the similarity between pretraining data and target data for known downstream tasks. Experiments show that our methods achieve results that are competitive with or even better to approaches using external models or datasets. Additionally, negCLIPLoss and NormSim are compatible with existing top techniques, allowing us to achieve a new state-of-the-art by combining them.
A notable limitation of our study is the exclusion of larger pretraining datasets, such as the large and xlarge scales of DataComp. However, DataComp-medium is the most commonly used benchmark for data selection in CLIP pretraining, and our method has demonstrated both effectiveness (Table 1-3) and efficiency (Table 4) on it. Future directions include exploring better ways to merge data selected by different methods and incorporating our methods into data scheduling scenarios.
Appendix A Theoretical Interpretation
@@ A.1 Theoretical Setup
For any ${\bm{x}}^{v},{\bm{x}}^{l}\in\mathbb{R}^{d}$ observable image and text training pairs, we define ${\bm{z}}^{v},{\bm{z}}^{l}$ to be the corresponding latent vectors which contain all semantically pertinent information about our tasks of interest. Similar to previous theoretical work [42], we assume each i.i.d pair ${\bm{z}}^{vl}$ follows zero-mean sub-gaussian distribution whose cross-covariance satisfies
@@
According to Lemma 4.1 in [42], using the CLIP loss to optimize the linear model has approximately the same training dynamics as using the regularized linear loss. Therefore, here we assume that we are learning $G_{v},G_{l}$ by maximizing the clip score gap between the contrastive pairs, plus a regularizer,
@@
A.2 Generalization Guarantees
We now provide theoretical guarantees and postpone our proof into Appendix A.3. Firstly, we are going to prove the intuition behind $\text{NormSim}_{2}$score.
Lemma A.1 (Intuition behind $\text{NormSim}_{2}$).

### v2

@@ 4 Experimental Results
In this section, we evaluate the performance of negCLIPLoss and NormSim, aiming to address the following questions: Q1: Given a fixed CLIP teacher model, can our methods more effectively utilize CLIP embeddings for data filtering? Q2: Are our methods applicable to diverse CLIP teacher models with varying architectures or different pretrained datasets? Q3: How does our method compare to other leading approaches that utilize external models or multimodal datasets? Could our method be compatible with these methods?
@@ 4.1 Setup
We adhere to the standardized training and evaluation protocols of the DataComp benchmark (gadre2023datacomp). We employ the medium-scale training configuration of DataComp (DataComp-medium). It provides a substantial dataset comprising 128 million low-quality, web-curated image-text pairs to be filtered. Once the data subset is obtained by some data filtering strategy, it will be used to train a fixed CLIP-B/32 model in a fixed training budget that allows the model to pass 128 million data points an epoch. Therefore, smaller subsets will be repeated more frequently, ensuring a fair comparison. We note that the size of the DataComp dataset becomes smaller over time since some URLs of images become invalid [^2] See https://github.com/mlfoundations/datacomp/issues/3 and similar issues in maharana2023d2. , and we only successfully downloaded about 110M data. Therefore, the results of baselines on the leaderboard do not apply to our datasets. So we reproduce top baselines on the leaderboard with their public UIDs of the selected data [^3] The results of MLM (wang2024finetuned) are from their paper.. More details about evaluation tasks and architecture are in Appendix D.1.
@@ 4.2 Baselines
We restate the three current research directions mentioned before based on how much external resources are employed: (D1) using OAI CLIP alone while optimizing embedding employment strategies, (D2) training and using a more advanced CLIP embedding model based on external data, and (D3) utilizing non-CLIP external models to aid data selection [^4] Note that D2 and D3 may also incorporate strategies from D1. For example, CLIPScore (D1) has been used in almost all the top methods. Therefore, we categorize baselines by the largest possible category they encompass.. In our experiment parts, we consider baselines following the above categorization: (1) D1, including CLIPScore (hessel2021clipscore), Image-based filtering (gadre2023datacomp), $\mathbb{D}^{2}$ Pruning (maharana2023d2). (2) D2, D3, including DFN (fang2023data), HYPE (kim2024hype), HYPE $\cup$ DFN (previous state-of-the-art), T-MARS (maini2023t), Devils (yu2023devil), and MLM (wang2024finetuned). Please refer to Fig. 4, Appendix C, and Appendix D.5 for more details. Intuitively, performance should be ranked as D2, D3 > D1. However, our results show that cross-setting comparisons are possible and our D1 methods can perform similar or even better than most of D3 methods.
Table 1: Results on DataComp-medium from methods that use only OpenAI’s CLIP-L/14 model, i.e., all methods are from the D1 category. The “dataset size” represents the size of the subset obtained from different approaches. NormSim(IN-1k) denotes using the training data of ImageNet-1k as the target while NormSim(Target) represents using that of all 24 available downstream tasks. Complete Results in Appendix E.2.
Filtering Strategy Dataset IN-1k IN Shift VTAB Retr. Avg.
@@
No filtering (gadre2023datacomp) 110M 17.3 15.0 25.2 21.3 25.6
CLIPScore (20%) (hessel2021clipscore) 22M 25.4 22.7 31.8 22.0 31.0
CLIPScore (30%) (hessel2021clipscore) 33M 26.4 23.6 32.6 24.5 32.2
Image-based (gadre2023datacomp) 24M 25.5 21.9 30.4 24.6 29.9
CLIPScore (30%) $\cap$ Image-based 11M 27.4 23.9 31.9 21.4 30.8
$\mathbb{D}^{2}$ Pruning (maharana2023d2) 22M 23.2 20.4 31.4 18.7 29.5
@@
4.3 Main Results and Discussions
4.3.1 D1 Category (Q1)
In Table 1, we compare D1 methods where only the OAI CLIP model is allowed to be used. We show that our Methods leverage OAI CLIP-L/14 better. First, negCLIPLoss outperforms CLIPScore on all metrics, regardless of whether it is used alone or combined with other methods. These results support our claim that negCLIPLoss can more accurately estimate the data quality. Second, when target knowledge is available, $\text{NormSim}_{\infty}$ can further improve filtering significantly. Especially, compared with the best baseline “CLIPScore (30%)”, our “negCLIPLoss $\cap$ $\text{NormSim}_{\infty}$(Target)” improves 5.3% on ImageNet-1k and 2.8% on average 38 downstream tasks, respectively. Later in Table 3 we will see that this result outperforms all the D3 baselines except DFN $\cup$ HYPE. And we can also see that using ImageNet-1k-only target data can still produce good results.
@@ 4.3.1 D1 Category (Q1)
4.3.2 Universality (Q2)
To evaluate whether our method applies to other CLIP teacher models, we replaced OAI-L/14 with OAI-B/32 and DFN-P as embedding models. Results are in Table 2 and Appendix E.3.555The original DFN paper selects a subset comprising 19.2M data points, which accounts for approximately $17.5\%$ of our dataset and $15\%$ of their dataset, we incorporate these sampling ratios into our comparison. We conclude that (1) negCLIPLoss can be applied to different CLIP embedding models. Our proposed negCLIPLoss can consistently outperform CLIPScore on all downstream tasks for different filtering ratios and models, like a 0.5%-5.4% increase on ImageNet-1k. (2) Embedding required by NormSim should have good downstream performance. When combining negCLIPLoss with $\text{NormSim}_{\infty}$, OAI-B/32 obtains results even better than those using OAI-L/14 (Table 1), while DFN-P achieves results even worse than using negCLIPLoss alone. The reason is that, unlike OAI CLIP, DFN-P is specially designed for data filtering at the expense of downstream task performance [^6] For example, the ImageNet-1k accuracy for DFN-P, OAI CLIP-B/32, and OAI CLIP-L/14 are 45%, 63%, and 75%, respectively.. This indicates that the embeddings obtained from DFN on target data might be highly unreliable, leading to inaccurate NormSim. Furthermore, if we use DFN-P for negCLIPLoss but OAI-B/32 for NormSim ("negCLIPLoss (17.5%) $\cap$ $\text{NormSim}_{\infty}^{\text{B/32}}$(Target)"), we can achieve the best results.
@@ 4.3.3 Comparison with D2 & D3 Categories (Q3)
In this part, we compare all the D2 & D3 baselines mentioned in Sec. 4.2 together with our best strategy in Table 3.  [^7] For “A $\cup$ B” mentioned in Table 3, we follow the way of Kim et al. (kim2024hype) to merge the data, which generates the subset separately for each method and then merge them. This will result in oversampling the shared data, which is intuitively more important. We claim: (1) Our methods can outperform most of the D3 methods. In Table 3, we show that without using any external models or data, our best combination, i.e., using OAI-B/32 for “negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)” (Ours (20%)), still outperforms all methods except DFN and “DFN $\cup$ HYPE”. This answers the first part of Q3 and further indicates that some external models may be redundant since CLIP embeddings already contain necessary information. (2) We can further improve the SOTA method. We can further boost the performance of the current SOTA method “HYPE $\cup$ DFN” by 0.9% on both ImageNet-1k and on average 38 downstream tasks, and close results can be achieved even without combining HYPE which utilizes the external embedding model MERU (desai2023hyperbolic). Here we use the data selected by both OAI-B/32 and L/14, which we found is more robust than using one of them alone. Our better results answer the second part of Q3, that is, our methods can be compatible with other D2&D3 methods.
Table 3: Results of all D1&D2&D3 top methods on DataComp-medium. “Ours (20%)” refers to use “negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)” to get 20% of original data, while “Ours (10%)” denotes applying “negCLIPLoss (20%) $\cap$ $\text{NormSim}_{\infty}$(Target)” to get 10%. And we use “*” to indicate the case where we choose the intersection of the data selected by using OAI CLIP-B/32 and OAI CLIP-L/14 separately, which results in about 15M data for “Ours (20%)*” and 7.4M data for “Ours (10%)*”.
Type Filtering Strategy Size IN-1k IN Shift VTAB Retr. Avg.
(1) (5) (11) (3) (38)
D2 DFN (fang2023data) 16M 36.0 30.1 36.2 27.0 35.4
D3 T-MARS (maini2023t) 22M 30.8 26.3 34.8 25.4 34.1
D3 Devil (yu2023devil) 20M 31.0 26.7 35.9 24.7 34.5
D3 MLM (wang2024finetuned) 38M 30.3 25.6 36.0 29.0 34.5
D3 HYPE (kim2024hype) 10M 30.3 25.8 34.3 22.2 31.9
D3 DFN $\cup$ HYPE (kim2024hype) 20M 36.4 30.8 38.5 28.0 36.8
@@
5 Conclusion and Reproducibility Statement
In this paper, we introduce negCLIPLoss and NormSim to enhance data selection in multimodal contrastive learning without relying on external resources. negCLIPLoss provides a more accurate quality metric compared to the commonly used CLIPScore, while NormSim measures the similarity between pretraining data and target data for known downstream tasks. Experiments show that our method demonstrates both effectiveness (Table 1-3) and efficiency (Table 4). Additionally, negCLIPLoss and NormSim are compatible with existing top techniques, allowing us to achieve a new state-of-the-art by combining them. All experiments are reproducible using code https://github.com/ypwang61/VAS/tree/negcliploss_normsim.
Appendix A Additional Discussion
A.1 Discussion about NormSim
Necessity of using vision-only information.
We use only the visual information $x^{v}$ instead of multi-modal information $x^{vl}$ for measuring similarity. This is because common crawled text often has brief captions, making the OAI CLIP language embedding weaker than its visual embedding model (gadre2023datacomp; shen2021much; zeng2022multi; yamada2022lemons). Consequently, the language part cannot characterize the pre-training and downstream task distribution as well as the visual part. This phenomenon is also observed in Gadre et al. (gadre2023datacomp), where image-based filtering (select data whose image embeddings are similar to that from ImageNet-1k) outperforms text-based filtering (select data whose captions contain words from ImageNet-21k). More ablation studies are provided in Appendix E.5.
Using proxy when downstream $X_{\text{target}}$ is inaccessible.
Surprisingly, we show that the 2-norm can also be used when only the pre-training set is available. In this case, we construct a proxy “target” set from the pre-training set itself. Specifically, let $S_{i}$ be the selected subset at step $i$, then we treat the current $S_{i}$ as the proxy “target” set. To construct the next smaller set, we greedily select the next data batch $S_{i+1}$ satisfying $\operatorname*{arg\,max}_{S_{i+1}\subset S_{i}}\sum_{x\in S}\text{{$\text{NormSim}_{2}$}}(S_{i},x),$ until reaching an N size subset. We call this approach $\text{{NormSim}}_{2}$-D (Dynamic) and will specify the algorithm details in Appendix D.4.
Theoretical justification.
Unlike many existing methods that force diversity by selecting training samples around each ${\bm{x}}_{\text{target}}$, our strategy greedily maximizes similarity without considering actual coverage. For the $p=2$ case, we demonstrate that maximizing $\text{NormSim}_{2}$ is optimal under a linear model $\bar{f}_{v}$, as shown in Appendix B. Our theorem also provides error guarantees for noisy embeddings and explains when vision-only embeddings outperform combined vision and language embeddings. Recent work by Joshi et al. (pmlr-v238-joshi24a) provides a similar analysis but focuses on high-quality data and cross-variance between images and texts. This approach is less effective than image-only methods for filtering noisy datasets, as discussed above.
Appendix B Theoretical Interpretation
@@ B.1 Theoretical Setup
For any ${\bm{x}}^{v},{\bm{x}}^{l}\in\mathbb{R}^{d}$ observable image and text training pairs, we define ${\bm{z}}^{v},{\bm{z}}^{l}$ to be the corresponding latent vectors which contain all semantically pertinent information about our tasks of interest. Similar to previous theoretical work (nakada2023understanding), we assume each i.i.d pair ${\bm{z}}^{vl}$ follows zero-mean sub-gaussian distribution whose cross-covariance satisfies
@@
According to Lemma 4.1 in (nakada2023understanding), using the CLIP loss to optimize the linear model has approximately the same training dynamics as using the regularized linear loss. Therefore, here we assume that we are learning $G_{v},G_{l}$ by maximizing the clip score gap between the contrastive pairs, plus a regularizer,
@@
B.2 Generalization Guarantees
We now provide theoretical guarantees and postpone our proof into Appendix B.3. Firstly, we are going to prove the intuition behind $\text{NormSim}_{2}$score.
Lemma B.1 (Intuition behind $\text{NormSim}_{2}$).